OpenStudy (anonymous):

Which conic section does this equation represent? x^2-4x+6y-9=5 A. Circle B. Ellipse C. Parabola D. Hyperbola

OpenStudy (anonymous):

parabola. C.

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

x^2 -4x +6y-9=5 (x-2)^2+6y+4-9=5 (x-2)^2+6y-5=5 (x-2)^2+6y=10 so parabola